# Odd-One-Out: Anomaly Detection by Comparing with Neighbors

### v1

@@ Abstract
This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances within a scene. Unlike the traditional AD benchmarks, in our setting, anomalies in this context are scene-specific, defined by the regular instances that make up the majority. Since object instances are often partly visible from a single viewpoint, our setting provides multiple views of each scene as input. To provide a testbed for future research in this task, we introduce two benchmarks, ToysAD-8K and PartsAD-15K. We propose a novel method that generates 3D object-centric representations for each instance and detects the anomalous ones through a cross-examination between the instances. We rigorously analyze our method quantitatively and qualitatively in the presented benchmarks.
@@ 1 Introduction
Anomaly detection (AD) [13, 34] aims to detect patterns that deviate from expected behavior. In standard computer vision AD benchmarks, the non-conforming (or anomalous) patterns in images are often either due to high-level variations such as introduction of an object instance from an unseen category [2, 8, 12] or low-level variations in object shape and texture [6, 16] (see Fig. 1(a)). In these benchmarks, the definitions of normal and anomalous patterns are typically not predefined but implicitly learned through representations and/or classifiers [21, 2, 16] that are invariant to appearance changes within the normal data distribution while being sensitive to the anomalous ones. However, in many real-world applications such as quality control in production lines, product specification, hence the definition of ‘normality’, is not only known in advance but also specific to each object instance. For example, while a coffee cup with a red handle is considered normal and one with a blue handle is anomalous when the goal is to produce red-handled cups, but the opposite is true if the goal is to produce blue-handled cups. This instance-specific definition of normality cannot be addressed by methods that rely on a global and implicit definition of normality.
In this paper, inspired from real-world visual inspection scenarios, we introduce a new AD problem, along with two new benchmarks with different characteristics and a novel solution. As depicted in Fig. 1 (b), in this proposed setting, each image contains a scene with multiple instances of the same object (e.g., coffee cups) including either only normal, or a mix of normal and anomalous instances where the anomalous instances constitute the minority. As defining precise visual specifications for shape and appearance is often infeasible, we assume that normal instances, which form the majority, provide a scene-specific reference for ‘normality’. Our objective is to detect anomalous instances in a scene while generalizing to previously unseen scenes including novel objects and spatial configurations.
@@
Unlike existing benchmarks, our task often requires a holistic understanding of the scene by comparing instances with each other, although some anomalies, such as cracks and misaligned parts, can be identified by inspecting individual instances. Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent, unlike the existing benchmarks that provide a single view as input (see Fig. 1(c)). Our goal is to detect anomalous samples in previously unseen scenes from multiple views.
Figure 1: The standard and our new AD settings.
@@
The proposed problem presents several challenges requiring the following capabilities: i) 3D understanding of the scene and registering the views from multiple camera viewpoints without groundtruth 3D knowledge while identifying potential occlusions, ii) aligning and comparing object instances with each other without their relative pose information in both training and evaluation, iii) learning representations that generalize to unseen object instances during testing. To address these challenges, we propose a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, produces a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances through an efficient attention mechanism. We leverage recent advances in differentiable rendering [31] and self-supervised learning [33] to supervise 3D representation learning. Specifically, we render the voxel representations for multiple viewpoints and match them with the given 2D views to produce geometrically consistent features. Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [33], enhancing the correspondence matching across instances. Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, including instances from common semantic object categories as well as mechanical parts respectively to provide a testbed for future research. Our method significantly outperforms various baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
@@ 2 Related Work
Few-shot AD. There are some works [18, 23, 52, 50, 7] aiming to detect anomalies from a small number of normal samples as support images. In our setting, the concept of normality in each image is also learned from a few instances only. However, unlike them, the concept of normality is scene-specific, and our method can generalize to previously unseen instances without requiring any modification by learning from a support set. In addition, our setting involves multi-object multi-view data samples as input, unlike the single-object single-view in theirs.
@@
Leveraging foundation models. Large-scale pretraining on image datasets has shown impressive generalization capabilities on various tasks [35, 25, 33]. Previous works [57, 3] demonstrate that features extracted from DINOv2 [33] serve as effective dense visual descriptors with localized semantic information for dense correspondence estimation task. Some recent works [26, 46] use feature distillation techniques to leverage 2D foundation vision models for 3D task. Inspired by these works we utilize DINOv2 to distill its dense semantic knowledge into our 3D network that enables our network to infer robust local correspondences, which aids in fine-grained object matching.
@@ 3 Method
3.1 Overview
Consider a scene containing multiple rigid objects $\{o_{n}\}_{n=1}^{N}$ of the same instance, where $N$ represents the number of objects in the scene. Each object’s pose is arbitrary and unknown. The goal is to identify anomalies in the group of objects by assessing their mutual similarity. An observation in our setting consists of $M$-view $H\times W$ dimensional RGB images =$\mathcal{I}=\{\bm{I}_{t}\}_{t=1}^{M}$, and their corresponding camera projection matrices $\mathcal{P}=\{\bm{P}_{t}\}_{t=1}^{M}$, $\bm{P}_{t}=\bm{K}\left[\bm{R}_{t}|\bm{T}_{t}\right]$, with intrinsic $\bm{K_{t}}$, rotation $\bm{R}_{t}$ and translation $\bm{T}_{t}$ matrices. In our setting, $M$ is $5$, forming sparse-view inputs. Our goal is to learn a mapping $\psi$ from the multi-view images to object-centric anomaly labels ${y}_{n}\in\{0,1\}$ and its corresponding 3D bounding box $\bm{b}_{n}$, defined as:
@@ 3.1 Overview
Note, labels $y_{n}$ are defined relative to other objects in the scene. For example, consider a group of three coffee cups, of which two have red handles, and one has a blue handle. In this example, the latter cup is considered an anomaly.
Figure 2: Overview of our framework. We extract features from a sequence of input views using a 2D CNN and then back-project them into a 3D volume, which is then refined using a 3D CNN, resulting in $\bm{F}_{v}$. We then extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$ using RoI pooling, where $\mathcal{B}$ estimates the RoI region of each object. The cross-instance matching module $\mathcal{M}$ then learns a correlation among the objects using a sparse voxel attention. To improve the 3D representation of the scene, we distill the knowledge of a 2D vision model namely DINOv2, and integrate the learned knowledge into our 3D network via differentiable rendering.
Our architecture, illustrated in Fig. 2, is comprised of three main components: First, the 3D feature fusion module encodes each view image and projects it to 3D, forming a fused 3D feature volume. Second, the feature distillation block is employed to enhance the 3D feature volume through differentiable rendering. This facilitates finding local correspondences between objects. Finally, the cross-instance matching module leverages established correspondences to compare all similar object regions in the scene using a sparse voxel attention mechanism. Next, we elaborate on details.
@@
We first extract 2D features $\bm{F}_{t}=\mathcal{E}_{2D}(\bm{I}_{t})\in\mathbb{R}^{d\times h\times w}$ for each input view using a shared CNN encoder $\mathcal{E}_{2D}$, where $d$ is the feature dimension. These 2D features are then projected into 3D voxel space as follows:
@@
where $\Pi_{\texttt{proj}}$ back-projects each view feature $\bm{F}_{t}$ using known camera intrinsic and extrinsic, generating 3D feature volumes of size $d\times v_{x}\times v_{y}\times v_{z}$. These feature volumes are aggregated over all input views using an average operation as in [32, 42]. Finally, a 3D CNN-based network $\mathcal{E}_{3D}$ is employed to refine the aggregated feature volume, resulting in a final voxel representation $\bm{F}_{v}$.
@@
$[\hat{\bm{I}}_{t},\hat{\bm{I}}_{t\sigma}]=\mathcal{R}([\bm{V}_{c},\bm{V}_{\sigma}],\bm{P}_{t}),\quad\mathcal{L}^{\text{im}}_{t}=||\bm{I}_{t}-\hat{\bm{I}}_{t}||^{2}+\lambda_{\sigma}||\bm{I}_{t\sigma}-\hat{\bm{I}}_{t\sigma}||^{2},$ (3)
where $\hat{\bm{I}}_{t}$ and $\hat{\bm{I}}_{t\sigma}$ are rendered image and mask. $\lambda_{\sigma}$ is a loss weight.
We propose to improve the 3D voxel representation $\bm{F}_{v}$ by also reconstructing neural features instead of just color and density. We supervise the feature reconstruction by a pretrained 2D image encoder $\Phi$ as a teacher network. We choose DINOv2 [33] as the teacher network due to its excellent ability to capture various object geometries and correspondences.
@@
We use a projector function $\beta$, implemented as a four-layered $1\times 1\times 1$ convolution block that projects $\bm{F}_{v}$ to a neural feature field $\bm{V}_{f}$, changing the channel dimension from $d$ to $d_{f}$. Similar to color rendering, we generate rendered features $\hat{\Phi}_{t}$ of size $d_{f}\times h_{f}\times w_{f}$ at a given viewpoint using volume renderer $\mathcal{R}$. The objective is to minimize the difference between the rendered features and the teacher’s features $\Phi(\bm{I}_{t})$. We choose cosine distance as our feature loss ($\mathcal{L}^{\text{feat}}_{t}$) which we find easier to optimize compared to the standard L2 loss. We apply stop-gradient to density in rendering of features $\hat{\Phi}_{t}$, as the teacher’s features are not fully multi-view consistent [3], which could harm the quality of reconstructed geometry. The final reconstruction loss is the sum of all image and feature reconstruction losses weighed by a factor $\lambda_{f}$:
$\mathcal{L}^{r}=\sum_{t=1}^{M}(\mathcal{L}^{\text{im}}_{t}+\lambda_{f}\mathcal{L}^{\text{feat}}_{t}).$ (4)
@@
The key benefits of reconstructing DINOv2 features are twofold. 1) Firstly, distilling features from general-purpose feature extractors pre-trained on large external datasets, incorporates open-world knowledge into the 3D representation. This enables our model to perform significantly better on unseen object instances or even on novel categories, as demonstrated in the experiments. 2) Secondly, the distillation enforces consistent 3D scene representation, leading to identical features for the same object geometries. This enables the model to infer robust local correspondences (see Fig. 4), which aids in fine-grained object matching.
3.3 Object-centric 3D Features Extraction
We extract bounding box regions of objects using a box estimator $\mathcal{B}$. First, we obtain a voxel reconstruction of the scene by applying a threshold to the predicted density $\bm{V}_{\sigma}$. Then, we employ DBScan [20], a density-based clustering method to retrieve all bounding box regions $\{\bm{b}_{n}\}_{n=1}^{N}$ corresponding to the objects in the scene. Next we obtain object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, each with a size of $8\times 8\times 8$ from these regions through applying RoI pooling [17].
To extract correspondence indices given two objects with feature volumes $\bm{z}_{n}$ and $\bm{z}_{m}$, We formally define a function $\mathcal{C}_{k}$:
$\mathcal{C}_{k}(\bm{z}_{n},\bm{z}_{m})=\texttt{top}_{k}[\beta(\bm{z}_{n})^{T}\beta(\bm{z}_{m})].$ (5)
The function returns top-$k$ most relevant feature locations in $\bm{z}_{m}$ for each voxel location in $\bm{z}_{n}$. This is achieved by first projecting each voxel feature using the same projector function $\beta$ and then computing their voxel-level pairwise similarity. We use $\mathcal{C}_{k}$ to perform sparse attention-based comparisons between multiple object volumes, as described next.
@@ 3.4 Cross-instance Matching
The cross-instance matching module $\mathcal{M}$ takes the object features $\{\bm{z}_{n}\}_{n=1}^{N}$, effectively learns to correlate them using sparse voxel attention, and subsequently predicts their object-specific labels.
Unlike the vanilla self-attention module in standard transformers [19] that uses all tokens for the attention computation, which is inefficient for our task and may introduce noisy interactions with irrelevant features, potentially degrading performance. To overcome this, we compute the sparse voxel attention only among geometrically corresponding voxel locations, as shown in the inset.
Let $z_{n}[i]\in\mathbb{R}^{d}$ denote the $i$-th voxel of the $n$-th object volume in the scene. The query, key, and value embeddings are calculated using linear projections as:
$\bm{Q}_{n}[i]=\bm{W}^{Q}{\bm{z}}_{n}[i],\quad\bm{K}_{n}[i]=\bm{W}^{K}{\bm{z}}_{n}[i],\quad\bm{V}_{n}[i]=\bm{W}^{V}{\bm{z}}_{n}[i],$ (6)
where the weights $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are shared across all objects.
Let $c_{k}^{nm}[i]$ denote the set of all corresponding voxel indices obtained using Eq. 5. Then, the attention is calculated as:
$\bar{\bm{z}}_{n}[i]=\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}\sum_{j\in c_{k}^{nm}[i]}\text{softmax}\left(\frac{\bm{Q}_{n}[i]\bm{K}_{m}[j]}{\sqrt{d}}\right)\bm{V}_{m}[j].$ (7)
The updated feature volume $\bar{\bm{z}}_{n}$ is passed through 3D CNN blocks to downsample by a factor of $1/8$, which is finally reshaped into a vector and fed to a 2-layer MLP outputting the final prediction $\hat{y}_{n}$. The classification loss is calculated as:
@@
where $\ell_{\text{bce}}$ is the binary cross-entropy loss function. The total training loss of our framework is $\mathcal{L}=\mathcal{L}^{\text{bce}}+\lambda_{p}\mathcal{L}^{r}$, where $\lambda_{p}$ is a loss weight. We employ stage-wise training: first pretraining with only the reconstruction loss ($\mathcal{L}^{r}$), followed by end-to-end training with both losses.

### v2

@@ Abstract
This paper introduces a novel anomaly detection (AD) problem that focuses on identifying ‘odd-looking’ objects relative to the other instances in a given scene. In contrast to the traditional AD benchmarks, anomalies in our task are scene-specific, defined by the regular instances that make up the majority. Since object instances may be only partly visible from a single viewpoint, our setting employs multiple views of each scene as input. To provide a testbed for future research in this task, we introduce two benchmarks, ToysAD-8K and PartsAD-15K. We propose a novel method that constructs 3D object-centric representations from multiple 2D views for each instance and detects the anomalous ones through a cross-instance comparison. We rigorously analyze our method quantitatively and qualitatively on the presented benchmarks. The datasets, source code, and models will be made available upon publication.
Figure 1: (a) We propose a new anomaly detection task focused on identifying ‘odd-looking’ objects relative to other instances within a scene. Inspired by real-world quality control in production environments, this task aims to detect subtle variations in geometry and texture, including defects like cracks and fractures, in a group of manufactured samples. (b) Our setting is scene-specific, requiring a comparison of object instances within the input scene, unlike the standard AD setting, which takes only a single object as input. (c) Our matching-based paradigm enables cross-category performance.
@@ 1 Introduction
Anomaly detection (AD) [10, 29] aims to detect patterns that deviate from expected behavior. In standard AD benchmark for computer vision, anomalies may arise from high-level factors such as introduction of an object instance from an unseen category [9, 31, 2] or low-level ones such as defects and variations in the normal object shape and texture [8, 6, 13]. In these benchmarks, AD methods analyze each object individually, ignoring relationships between objects in the same scene or contextual factors. However, in many real-world scenarios, ‘normality’ can be conditional. For instance, in a production line for blue-handled coffee cups, a yellow handle is anomalous, and in a production line for yellow-handled coffee cups, a blue handle is anomalous (see Fig. 1(b)). This context dependent normality cannot be addressed by methods assuming a fixed standard of normality.
@@
Inspired by real-world visual inspection, we introduce a new AD problem with two new benchmarks and a novel solution. In our setup (see Fig. 1(a)), each image contains a scene with multiple instances of the same object (e.g., coffee cup) including either all normal, or a mix of normal and anomalous instances. Since defining exact shape and appearance specifications on paper is often impractical, we assume each scene has multiple normal instances together providing a scene-specific reference for ‘normality’. Our objective is to detect anomalous instances in a scene and to generalize to previously unseen scenes including unseen objects and layouts. Unlike existing benchmarks, this task often requires a holistic understanding of the scene by comparing instances with each other with the exception of few anomalies such as cracks and misaligned parts that can be identified by inspecting individual instances. Additionally, since performing AD from a single view can be ambiguous due to self-occlusion and occlusion between the instances, we provide multiple views of the scene to cover its entire relevant extent unlike existing benchmarks with single-view inputs (see Fig. 1(a)).
The proposed problem presents several challenges requiring the following key capabilities: i) achieving 3D understanding of the scene and view registration from multiple camera viewpoints without groundtruth 3D knowledge while handling occlusions, ii) aligning and comparing object instances without their pose information during training and testing, iii) learning representations that generalize to unseen object instances from seen or unseen categories at test time.
@@
To tackle these challenges, we introduce a novel method that takes as input multiple views of the same scene, projects them into a 3D voxel grid, generates a 3D object-centric representation for each instance, and predicts their labels by cross-correlating instances using an efficient attention mechanism. We utilize advances in differentiable rendering [26] and self-supervised learning [28] to supervise 3D representation learning. Specifically, we render the voxel representations for multiple viewpoints and match them with the given 2D views to produce geometrically consistent features. Additionally, we encourage our model to learn part-aware 3D representations by distilling features from 2D self-supervised model DINOv2 [28], enhancing the correspondence matching across instances and their alignment. Finally, since there is no prior benchmark for this task, we propose two new benchmarks – ToysAD-8K and PartsAD-15K, which include common objects categories as well as mechanical parts respectively to facilitate further research. Our method significantly outperforms baselines that do not compare instances with each other, and we rigorously analyze various aspects of our model and benchmarks.
@@ 2 Related Work
Few-shot AD. There are some works [15, 19, 46, 44] aiming to detect anomalies from a small number of normal samples as support images. In our setting, the concept of normality in each image is also learned from a few instances only. However, unlike them, the concept of normality is scene-specific, and our method can generalize to previously unseen instances without requiring any modification by learning from a support set. In addition, our setting involves multi-object multi-view data samples as input, unlike the single-object single-view in theirs.
@@
Figure 2: Overview of our framework. We extract features from a sequence of input views using a 2D CNN and back-project them into a 3D volume, which is then refined with a 3D CNN, resulting in $\bm{F}_{v}$. Next, we extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, which are fed into the cross-instance matching module to learn correlations among objects using sparse voxel attention. To improve the 3D representation of the scene, we distill the knowledge of a 2D vision model namely DINOv2, and integrate the learned knowledge into our 3D network via differentiable rendering. This aids in obtaining a part-aware and geometrically consistent 3D feature representation.
Leveraging large-scale models. Large-scale pretraining on image datasets has shown impressive generalization capabilities on various tasks [30, 21, 28]. Previous works [50, 3] demonstrate that features extracted from DINOv2 [28] serve as effective dense visual descriptors with localized semantic information for dense correspondence estimation tasks. Some recent works [22, 40] use feature distillation techniques to leverage 2D foundation vision models for 3D tasks. Inspired by these works, we utilize DINOv2 to distill its dense semantic knowledge into our 3D network enabling our network to infer robust local correspondences, which aids in fine-grained object matching.
@@ 3 Method
Consider a scene containing $N$ instances $\{o_{n}\}_{n=1}^{N}$ of the same rigid object where their poses are unknown. The goal is to identify anomalies in the group of objects by assessing their mutual similarity. An observation in our setting consists of $M$-view $H\times W$ dimensional RGB images =$\mathcal{I}=\{\bm{I}_{t}\}_{t=1}^{M}$, and their corresponding camera projection matrices $\mathcal{P}=\{\bm{P}_{t}\}_{t=1}^{M}$ where $\bm{P}_{t}=\bm{K}\left[\bm{R}_{t}|\bm{T}_{t}\right]$, with intrinsic $\bm{K_{t}}$, rotation $\bm{R}_{t}$ and translation $\bm{T}_{t}$ matrices. We set $M$ to $5$, forming sparse-view inputs in our experiments. Our goal is to learn a mapping $\psi$ from the multi-view images to object-centric anomaly labels ${y}_{n}\in\{0,1\}$ and its corresponding 3D bounding box $\bm{b}_{n}$, defined as:
@@
Note that labels $y_{n}$ are defined relative to other objects in the scene. For example, consider a group of three coffee cups, of which two have yellow handles, and one has a blue handle. In this example, the latter cup is considered an anomaly.
Our pipeline, illustrated in Fig. 2, is comprised of four main components: First, the 3D feature volume construction module (Sec. 3.1) encodes each view image and projects it to 3D, forming a fused 3D feature volume. Second, the feature enhancement block (Sec. 3.2), employed only during training, facilitates the learning of a part-aware and geometrically consistent 3D space through differentiable rendering and feature distillation. Third, the object-centric feature extraction module (Sec. 3.3) locates each object in the scene and represents it with a feature volume. Finally, the cross-instance matching module (Sec. 3.4) efficiently compares all similar object regions using a sparse voxel attention mechanism and predicts instance anomalous labels and 3D coordinates. Next, we elaborate on details.
@@
In our first module, we obtain a 3D feature volume representation given multi-view images as input. Specifically, we first extract 2D features $\bm{F}_{t}=\mathcal{E}_{2D}(\bm{I}_{t})\in\mathbb{R}^{d\times h\times w}$ for each input view using a shared CNN encoder $\mathcal{E}_{2D}$, where $d$ is the feature dimension. These 2D features are then projected into 3D voxel space as follows:
@@
where $\Pi_{\texttt{proj}}$ back-projects each view feature $\bm{F}_{t}$ using known camera intrinsic and extrinsic parameters, generating 3D feature volumes of size $d\times v_{x}\times v_{y}\times v_{z}$. These feature volumes are aggregated over all input views using an average operation as in [27, 37]. Finally, a 3D CNN-based network $\mathcal{E}_{3D}$ is employed to refine the aggregated feature volume, resulting in a final voxel representation $\bm{F}_{v}$ of size $d\times v_{x}\times v_{y}\times v_{z}$, which is further fed into the feature enhancement block, described next, to obtain a part-aware and geometrically consistent 3D representation.
3.2 Feature Enhancement Block
@@
$\begin{split}\mathcal{L}^{\text{im}}_{t}=||\bm{I}_{t}-\hat{\bm{I}}_{t}||^{2}+||\bm{I}_{t\sigma}-\hat{\bm{I}}_{t\sigma}||^{2},\end{split}$ (3)
where $[\hat{\bm{I}}_{t},\hat{\bm{I}}_{t\sigma}]=\mathcal{R}([\bm{V}_{c},\bm{V}_{\sigma}],\bm{P}_{t})$, $\hat{\bm{I}}_{t}$ and $\hat{\bm{I}}_{t\sigma}$ are the rendered image and mask.
@@
Moreover, we improve the 3D voxel representation $\bm{F}_{v}$ by reconstructing neural features in addition to geometry and appearance. We supervise the feature reconstruction by a pretrained 2D image encoder $\Phi$ as a teacher network. We choose DINOv2 [28] as the teacher network due to its excellent ability to capture various object geometries and correspondences. To this end, we use a projector function $\beta$, instantiated as a four-layered $1\times 1\times 1$ convolution block that projects $\bm{F}_{v}$ to a neural feature field $\bm{V}_{f}$, changing the channel dimension from $d$ to $d_{f}$. Similar to color rendering, taking $\bm{V}_{f}$ and $\bm{V}_{\sigma}$ as input, we generate rendered features $\hat{\Phi}_{t}$ of size $d_{f}\times h_{f}\times w_{f}$ at a given viewpoint using volume renderer $\mathcal{R}$, and minimize the difference between the rendered features and the teacher’s features $\Phi(\bm{I}_{t})$. We choose cosine distance as our feature loss ($\mathcal{L}^{\text{feat}}_{t}$), which we find easier to optimize compared to the standard L2 loss. We do not allow the gradients to follow through the density in the rendering of features $\hat{\Phi}_{t}$, as the teacher’s features are not fully multi-view consistent [3], which could harm the quality of reconstructed geometry. The final reconstruction loss is the sum of all image and feature reconstruction losses:
$\mathcal{L}^{r}=\sum_{t=1}^{M}(\mathcal{L}^{\text{im}}_{t}+\mathcal{L}^{\text{feat}}_{t}).$ (4)
@@
The key benefits of reconstructing DINOv2 features are twofold. First, distilling features from general-purpose feature extractors pretrained on large external datasets incorporates open-world knowledge into the 3D representation. This enables our model to perform significantly better on unseen object instances or even on novel categories, as demonstrated in the experiments. Second, the distillation through rendering enforces consistent 3D scene representation, leading to very similar features for the same object geometries. This enables the model to infer robust local correspondences (see Fig. 6), which aids in fine-grained object matching.
3.3 Object-centric Feature Extraction
Having the enhanced 3D feature volume representation, we aim to localize each object instance in the scene and represent each with a 3D feature volume such that the instances can be compared to each other in the next step. To this end, we use the predicted density volume $\bm{V}_{\sigma}$, obtained in Sec. 3.2, to extract a coarse point-cloud structure of the scene by applying a threshold to $\bm{V}_{\sigma}$. Then, we employ DBScan [17], a density-based clustering method to segment the foreground objects in the point cloud. This produces bounding box regions $\{\bm{b}_{n}\}_{n=1}^{N}$ corresponding to the objects in the scene. Finally, we apply RoI pooling [14] on the voxel representation $\bm{F}_{v}$ to extract object-centric feature volumes $\{\bm{z}_{n}\}_{n=1}^{N}$, each with a size of $d\times 8\times 8\times 8$. Note that we do not require ground truth box coordinates during training. Next, we pass them to the cross-instance matching module for fine-grained comparison.
@@ 3.4 Cross-instance Matching
Our task requires comparing object instances with each other, which further demands geometric alignment for each pair. Since the relative pose of each instance with respect to the camera is unknown, we avoid explicit alignment and instead focus on finding local matches between them. To this end, for a given pair of object instances ($m$ and $n$), we compute the most similar local features between their feature volumes ($\bm{z}_{n}$ and $\bm{z}_{m}$) respectively:
$\mathcal{C}^{nm}_{k}[i]=\texttt{top}_{k}[\beta(\bm{z}_{n}[i])^{T}\beta(\bm{z}_{m}[.])],$ (5)
where we denote the operation as $\mathcal{C}^{nm}_{k}$, which returns top-$k$ most relevant feature locations in $\bm{z}_{m}$ for the $i$-th voxel location in $\bm{z}_{n}$. This is achieved by first projecting each voxel feature utilizing the projector function $\beta$ (described in Sec. 3.2 and then computing their voxel-level pairwise similarity followed by a $\texttt{top}_{k}$ operation. We use $\mathcal{C}^{nm}_{k}$ to perform sparse attention-based comparisons between multiple object volumes, as described next.
The query, key, and value embeddings are calculated using linear projections as:
$\begin{split}\bm{Q}_{n}[i]=\bm{W}^{Q}{\bm{z}}_{n}[i],\quad\bm{K}_{n}[i]=\bm{W}^{K}{\bm{z}}_{n}[i],\\ \quad\bm{V}_{n}[i]=\bm{W}^{V}{\bm{z}}_{n}[i],\end{split}$ (6)
where the weights $\bm{W}^{Q}$, $\bm{W}^{K}$ and $\bm{W}^{V}$ are shared across all objects. Then, the attention is calculated as:
$\bar{\bm{z}}_{n}[i]=\sum_{\begin{subarray}{c}m=1\\ m\neq n\end{subarray}}^{N}\sum_{j\in\mathcal{C}_{k}^{nm}[i]}\text{softmax}\left(\frac{\bm{Q}_{n}[i]\bm{K}_{m}[j]}{\sqrt{d}}\right)\bm{V}_{m}[j].$ (7)
Unlike the vanilla self-attention module in standard transformers [16] that uses all tokens for the attention computation, which is inefficient for our task and may introduce noisy interactions with irrelevant features, potentially degrading performance. To overcome this, we compute the sparse voxel attention only among geometrically corresponding voxel locations using $\mathcal{C}_{k}^{nm}$.
Next, the updated feature volume $\bar{\bm{z}}_{n}$ is passed through 3D CNN blocks to downsample by a factor of $1/8$, which is finally reshaped into a vector and fed to a 2-layer MLP outputting the final prediction $\hat{y}_{n}$. The classification loss is calculated as:
@@
where $\ell_{\text{bce}}$ is the binary cross-entropy loss function. The total training loss of our framework is $\mathcal{L}=\mathcal{L}^{\text{bce}}+\mathcal{L}^{r}$.